Use the method of elimintion to solve system
2x+y+3z=30
4x-3y+6z=50
x+y-2z=30

Answer by stanbon(75887)

Use the method of elimintion to solve system
2x+y+3z=30
4x-3y+6z=50
x+y-2z=30
------
List the 3rd equation 1st:
x+y-2z=30
2x+y+3z=30
4x-3y+6z=50
------
Subtract 2 times the 1st from the 2nd
Subtract 4 times the 1st from the 3rd
-------
x+y-2z=30
0-y+7z=-30
0-7y+14z=-70
--------
Subtract 7 times the 2nd to the 3rd:
x + y - 2z = 30
0 - y + 7z = -30
0 + 0 -35z = 140
--------
Divide both sides by -35 to get:
x + y - 2z = 30
0 - y + 7z = -30
0 + 0 + z = -4
------
Solve the 2nd equation for "y":
-y -28 = -30
y = 2
-----
Solve the 1st equation for "x":
x + 2 + 8 = 30
x = 20
----
solution: x = 20; y = 2; z = -4
====================================
Cheers,
Stan H,
--------